### v1

# Deontic Temporal Logic for Formal Verification of AI Ethics

Abstract
Ensuring ethical behavior in Artificial Intelligence (AI) systems amidst their increasing ubiquity and influence is a major concern the world over. The use of formal methods in AI ethics is a possible crucial approach for specifying and verifying the ethical behavior of AI systems. This paper proposes a formalization based on deontic logic to define and evaluate the ethical behavior of AI systems, focusing on system-level specifications, contributing to this important goal. It introduces axioms and theorems to capture ethical requirements related to fairness and explainability. The formalization incorporates temporal operators to reason about the ethical behavior of AI systems over time. The authors evaluate the effectiveness of this formalization by assessing the ethics of the real-world COMPAS and loan prediction AI systems. Various ethical properties of the COMPAS and loan prediction systems are encoded using deontic logical formulas, allowing the use of an automated theorem prover to verify whether these systems satisfy the defined properties. The formal verification reveals that both systems fail to fulfill certain key ethical properties related to fairness and non-discrimination, demonstrating the effectiveness of the proposed formalization in identifying potential ethical issues in real-world AI applications.
Index Terms:
Artificial Intelligence, Ethics, Deontic Temporal Logic
I Introduction
Artificial Intelligence (AI) systems are becoming increasingly ubiquitous and influential in our lives, making decisions that can have significant ethical implications. As AI continues to advance and take on more complex tasks, it is crucial to ensure that these systems behave ethically [1, 2, 3, 4, 5, 6, 7, 8, 9]. However, defining and enforcing ethical behavior in AI is a challenging task, as ethics often involve abstract concepts and context-dependent judgments[10, 11, 12]. There are numerous principles generated by various organizations and regulation bodies. For instance, the Ethically Aligned Design (EAD) guidelines of IEEE recommend that AI design prioritize maximizing benefits to humanity [13]. Furthermore, The European Commission has released Ethics Guidelines for Trustworthy AI, stressing the importance of AI being human-centric [14]. The national plan for AI in the United Kingdom suggests the establishment of an AI Code [15]. Australia has also introduced its AI ethics framework [16], which adopts a case study approach to examine fundamental ethical principles for AI and offers a toolkit for integrating ethical considerations into AI development. Adding to this are Beijing’s AI principles, Amnesty International ACM code of ethics, and many more. In addition to governmental organizations, prominent companies such as Google [17] and SAP [18] have publicly released their AI principles and guidelines. Moreover, professional associations and non-profit organizations like the Association for Computing Machinery (ACM) have issued their recommendations for ethical and responsible AI [19, 20].
Despite these efforts, a consensus on the ethics of AI remains challenging. They lack a unified framework of guidelines that can be universally adopted by organizations, governments, and regulatory bodies to formulate and assess the ethics of systems. It is not yet clear what common principles and values AI should adhere to. Establishing cohesive and widely accepted ethical principles for AI is crucial across different organizations and domains. Moreover, ethics is a philosophical question of what is right or wrong [21, 22]. Its qualitative nature makes it complex and hard to define precisely and hence needs a mathematically rigorous framework.
To address this challenge, we are exploring the use of formal methods to express and prove the ethical correctness of AI systems. One promising approach is the use of deontic logic, a branch of modal logic that deals with concepts such as obligation, permission, and prohibition [23, 24]. Deontic logic provides a rigorous framework for reasoning about ethical norms and can be used to formalize ethical principles [25] and constraints. Several works have explored using deontic logic to formalize machine ethics mainly for robots [26, 27] and normative systems [28]. These studies have concentrated on Kantian ethics, integrating deontic and temporal logic to verify the ethical behavior of autonomous systems, such as unmanned aircraft, over time [29, 30].
While promising, these methods are often constrained by specific ethical frameworks and fail to scale with modern AI’s complexity, which increasingly mimics human tasks, leverages natural language processing, and operates on vast datasets. This leads to a proliferation of potentially subjective ethical rules influenced by personal biases. The dynamic, evolving nature of AI further complicates ethical formalizations [31]. Critically, many of these approaches remain theoretical, lacking practical integration with machine learning techniques, highlighting the need for more adaptive and implementable ethical frameworks in AI [32]. Our work represents a foundational effort to develop a unified framework that addresses ethical principles in AI systems, specifically focusing on granular levels of explainability and fairness. It builds upon existing approaches [33, 26] in specialized domains, extending them to tackle the unique challenges posed by modern AI ethics. In doing so, this paper introduces a novel direction for formalizing and verifying the ethical principles of AI systems.
While the proposed framework is broadly applicable to the ethics verification of autonomous systems, this work specifically focuses on its application to AI. The scope of this work is to provide a conceptual foundation and framework for formalizing AI ethics using deontic and temporal logic. Rather than focusing on individual actions or decisions, our approach emphasizes system-level specifications. It involves defining ethical properties that AI systems should ideally meet. For example, an AI system that uses gender as a feature should avoid making decisions explicitly based on it. These properties ensure that systems are designed to identify and mitigate biases and ethical violations effectively. Defining properties like "forbidden to consider sensitive features in predictions," provides a way for analysis of commonly discussed properties of AI in a unified manner. This abstraction also helps to reduces the difficulty in formalizing each action or each type of user [31]. Additional details on formalizing the properties of AI systems and their verification are provided in Section IV.
The basic model for applying deontic logic to AI ethics uses first-order logic to define predicates (Table I) and axioms that capture ethical requirements. This model introduces variables and predicates such as $\mathit{x}$ to indicate an AI system, $\mathit{x}_{a}$ to indicate that system $\mathit{x}$ performs an action $\mathit{a}$, and $\mathcal{E}(x)$ to indicate ethical behavior. Axioms 1.1,  1.2,  1.3, and  1.4 are then defined using these predicates to express ethical obligations, prohibitions, and permissions for AI systems. Building upon the basic model, an extended version incorporates temporal operators from temporal logic to reason about the ethical behavior of AI systems over time. This extension allows for the expression of more complex ethical requirements, such as the obligation for AI systems to maintain fairness over time or the prohibition of exhibiting bias. The temporal operators used in this model include “always” ($\Box$), “eventually” ($\diamond$), and “until” ($\mathcal{U}$) presented by Manna and Pnueli [34].
Theorems III.2 and III.3 in the basic model of deontic logic for AI ethics explore the relationships between ethical obligations, prohibitions, and permissions for AI systems. These theorems employ first-order logic and the defined predicates to derive conclusions about the ethical behavior of AI systems. The proofs of these theorems rely on techniques such as modus ponens, contraposition, and proof by contradiction to establish the logical connections between the axioms and the derived statements. The general flavor of the theorems is to provide a rigorous foundation for reasoning about the ethical requirements of AI systems, demonstrating how the Axioms 1.11.21.3, and  1.4 can be used to infer specific obligations, prohibitions, and permissions in various contexts. By establishing these logical relationships, the theorems contribute to a comprehensive framework for analyzing and ensuring the ethical behavior of AI systems. Similar to this, the Theorems III.5 to III.14 in the extended model, which incorporates temporal logic operators, explore the ethical behavior of AI systems over time. These theorems focus on capturing the temporal aspects of ethical requirements, such as the obligation to maintain fairness or the prohibition of exhibiting bias. The proofs of these theorems utilize the semantics of the temporal operators, such as $\Box$(always), $\diamond$(eventually), and $\mathcal{U}$(until), in conjunction with the Axiom lists 2 to 3 and predicates defined (Table I) in the basic model. The general flavor of the theorems in the extended model is to provide a more expressive and nuanced framework for reasoning about the ethical behavior of AI systems, considering the dynamic and evolving nature of these systems. The theorems establish logical connections between the temporal properties of AI systems and their ethical obligations, allowing for the analysis of more complex and realistic scenarios. By incorporating temporal aspects, the extended model enables a deeper understanding of the long-term ethical implications of AI systems and provides a foundation for designing and verifying AI systems that behave ethically over time.
The importance of this work lies in formalizing the axioms to define an AI system’s ethical requirements and its potential to provide a formal and verifiable framework for ensuring the ethical behavior of AI systems. Our experimental findings show the effectiveness of this formalization in assessing the ethics of real-world AI systems—Loan prediction and COMPAS. We evaluated the ethical aspects of the systems, wherein we defined specific properties that these systems must adhere to in order to be deemed ethical. Our results revealed that certain properties were indeed satisfied by the system, while others were not (Table III). The results demonstrated that applying deontic logic and temporal operators to AI ethics represents a significant step forward to formally specify and verify the ethical behavior of AI systems.
Section II discusses the related works. Section III introduces deontic logic for AI ethics formalization. Subsections III-B and III-C address fairness and explainability principles. Section IV covers the application of this approach to real-world AI systems, including algorithms 1 and 2 to demonstrate the implementation of this method on real datasets, providing readers with detailed insights into how it is executed. Section V concludes.
II Related Works
The field of ethical reasoning encompasses a range of approaches, often grounded in formal logic, to ensure trustworthy and morally sound behavior in autonomous systems. Several works contribute to this domain, presenting unique methodologies and frameworks to address the complex interplay between ethics and machine decision-making. Among these, deontological ethics, particularly Kantian frameworks, are well-suited for machine ethics due to their rule-based nature. This method ensures that machines refrain from harmful actions through rule-based formalization [35, 36, 33].
The earlier work introduce the GenEth ethical dilemma analyzer [37], which utilizes inductive logic programming to infer principles for ethical actions. Dominance Act Utilitarianism (DAU), a deontic logic of agency, is another framework for encoding and analyzing obligations in autonomous systems. DAU frameworks are efficient in addressing safety-critical behaviors, such as adherence to traffic laws and avoidance of reckless actions [38]. Such frameworks can formalize ethical obligations in systems like self-driving cars, enabling systematic reasoning about social and moral responsibilities [39]. Additionally, several works employ the Belief-Desire-Intention (BDI) framework to formalize reasoning about moral agents [40, 41, 30]. This structure supports transparency and formal verification in ethical decision-making processes for robots.
Further, the literature explores the use of high-level action languages and Answer Set Programming to design ethical autonomous agents [42, 43, 44]. There are several works that propose using deontic logic to constrain robot behavior in ethically sensitive environments, as this type of logic helps interpret natural language directly [45, 26, 27]. This frameworks are also used for ethical reasoning in the healthcare domain, emphasizing accountability and transparency [46]. Additionally, various works focus on using deontic logic-based frameworks for formalizing ethical reasoning in AI systems [47, 48, 49].
To accommodate the dynamic nature of machine environments [50], several studies propose integrating deontic logic with temporal operators, facilitating the representation of concepts like refraining from specific actions or opting for alternative actions [51, 52]. This extension facilitates a richer understanding of ethical constraints in dynamic environments. Furthermore, frameworks combining linear temporal logic with lexicographic preference modeling support ethical decision-making in robotics [32]. This literature survey provides a focused overview to contextualize the study, acknowledging the potential existence of other relevant works in the field.
Therefore, rule-based ethical theories, particularly deontology, are essential for developing trustworthy AI systems [53]. However, considering the dynamic nature of AI, especially regarding fairness and explainability at a granular level, significant gaps remain. Our work addresses these gaps by introducing fairness and explainability at multiple granularities, including stable, transient, inherent, and retrofitted/artificial dimensions. These distinctions capture the evolving nature of AI and its complex decision-making processes, providing a more comprehensive approach to ethical verification. Furthermore, existing frameworks often overlook the impact of personal biases introduced during training and lack mechanisms to mitigate them effectively. To address this, we propose an iterative learning approach designed to identify and reduce the influence of personal biases in the system. While prior research highlights the gap between theoretical ethical reasoning and its practical application in autonomous agents [32], our framework bridges this divide. By implementing and testing the framework in real-world AI systems such as COMPAS and Loan prediction systems, we validate its effectiveness and ensure its applicability.
A key feature of our approach is the generation of counterexamples that illustrate how specific properties may violate system specifications. This not only strengthens the verification process but also provides actionable insights for refining system behavior. Furthermore, we leverage theorem provers to capture and validate properties derived from real-world data distributions and predictions, ensuring alignment with ethical principles under varying conditions. By combining theoretical rigor with practical implementation, our framework advances existing methodologies, offering significant improvements in fairness, explainability, bias mitigation, and system-level validation. It establishes a structured and scalable approach for analyzing and verifying the ethical considerations of AI systems, setting a foundation for future research and development in ethical AI.
III Deontic Logic for Ethics
III-A Preliminaries
Deontic logic is a branch of symbolic logic that deals with normative concepts such as obligation ($O$), permission ($P$), and forbidden ($F$). Our work provides the reader with insight into the use of Deontic Logic to formalize and verify the ethical principles of an AI system. The principles that we focus in this work include fairness and explainability. The ethics of AI is more a philosophical question about what is morally right or wrong, permissible or impermissible. By representing ethical rules as deontic statements, AI designers can specify what a system ought or ought not to do. They can evaluate actions or decisions against a set of predefined ethical rules and determine whether the system complies with these rules. This is essential to guarantee that AI systems act morally following societal norms.
Standard Deontic Logic (SDL) and Temporal Deontic Logic (TDL) represent two distinct variations within deontic logic. We incorporate both SDL and its extension, TDL, into our work. SDL formulas include classical propositional logic and it operates as a monadic deontic logic, meaning its operators (obligation, permission, forbidden) apply to individual formulas ($\varphi$); they are read as "it is obligatory that $\varphi$", "it is permissible that $\varphi$", and "it is forbidden that $\varphi$" respectively. Furthermore, they are cross-definable. For instance, $P\varphi:=\neg O(\neg\varphi)$, and $F\varphi:=O\neg\varphi$ This logical statement explains that permission ($P$) or forbidden ($F$) can be represented in terms of obligation ($O$). Temporal Deontic Logic expands SDL by integrating temporal aspects into norms and obligations, introducing operators such as $always(\Box),eventaully(\diamond),next$, and $until(\mathcal{U})$. For instance, $next\ \varphi$ means that the proposition $\varphi$ holds in the next time step. Similarly, $\varphi\mathcal{U}\psi$ means $\varphi$ is true until $\psi$ becomes true. We use the semantics of the combined logic as an extension of the Kripke-style possible world semantics of deontic logic with temporal operators, as described in reference [54]. We recommend interested readers to refer [55, 54] for further details on the foundational principles.
In this section, we focus on formalizing the overall ethical behavior of an AI system. For the formalization, we use the predicates as shown in Table I. The predicate is a function that takes an input and returns a truth value. The following set of axioms has been formulated to articulate the necessary and sufficient characteristics for an AI system to be considered ethical.
Definition III.1 (TDL syntax).
Given a set $p$ of atomic proposition, the temporal deontic logic, TDL is defined as,
$\phi:=p|O\phi|P\phi|\neg P\phi|\Box\phi|\diamond\phi|\phi\mathcal{U}\psi$
Axiom 1.
Basic Model
  1. 1.1
    If an action $\mathit{a}$ is ethically required, then it is obligatory that an ethical AI system $\mathit{x}$ performs $\mathit{a}$: $\mathcal{E}(x)\rightarrow O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$.
  2. 1.2
    If an action $\mathit{a}$ is ethically prohibited, then it is forbidden for AI system $x$ to perform $\mathit{a}$: $\mathcal{E}(x)\rightarrow\neg P(\mathit{x_{a}}\land\neg\mathcal{E}(a))$.
  3. 1.3
    If an action $\mathit{a}$ is ethically permissible, then any AI system can perform $\mathit{a}$: $\mathcal{E}(x)\rightarrow P(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$.
  4. 1.4
    An AI system following ethical guidelines performs ethical actions: $\mathcal{G}(x)\rightarrow\mathcal{E}(a)$
TABLE I: Predicates used in this work
Predicate Explanation
$\mathcal{E}(x)$ $x$ exhibits ethical behavior
$\mathcal{G}(x)$ $x$ follows ethical guidelines
$\mathcal{F}(x)$ $x$ exhibits fairness
$\mathcal{B}(x)$ $x$ exhibits bias
$\mathcal{L}(x)$ $x$ learns iteratively
$\mathcal{X}(x)$ $x$ has inherent explainability
$\mathcal{R}(x)$ $x$ has retrofit explainability
$\mathcal{C}(x,c)$ $x$ is counterfactually fair given constraint $c$
$\mathcal{T}(x)$ $x$ exhibits transparency
$\mathcal{E}(a)$ An action $a$ is ethically required
Based on the foundational principles outlined above and the domain knowledge, we have formulated Theorems  III.2 and  III.3 to ensure that the AI system adheres to ethical standards in all relevant dimensions. Theorem III.2 asserts that for an AI system to be ethical, it must follow all the ethical guidelines. It specifies that the system should not only be accurate in predicting outcomes but also do so in a manner that upholds principles such as fairness, transparency, and non-discrimination.
Theorem III.2.
An AI system obliged to perform ethical action $\mathit{a}$ is obliged to follow ethical guidelines: $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))\rightarrow O(\mathit{x_{a}}\land\mathcal{G}(\mathit{x}))$
Proof.
Assume $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$ (AI system is obliged to perform ethical action).
[…]
Theorem III.3.
An ethical AI system is not permitted to refrain from an ethically required action $\mathit{a}$: $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))\rightarrow P(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$
Proof.
Assume $O(\mathit{x_{a}}\land\mathcal{E}(\mathit{a}))$ (the action is ethically required).
By contraposition, if it is ethically permissible for the AI system to refrain from the action, then the action is not ethically required:
[…]
However, this contradicts our assumption that the action is ethically required. Therefore, our initial assumption that it is ethically permissible for the AI system to refrain from the action must be false. ∎
While Theorems  III.2 and  III.3 contribute to formalizing the general ethics of an AI system, it remains essential to develop rigorous formalizations for each individual ethical principle.
III-B Formalizing Fairness
Let’s begin by formalizing the concept of fairness in an AI system by considering various scenarios where it must maintain fairness and where it might compromise it. It is an important aspect of an ethical AI and can be categorized into—transient fairness, and stable fairness. Existing literature suggests that a fair AI system should avoid considering the sensitive attributes of individuals in its decision-making process (Definition III.4). These attributes can potentially harm their sentiments, and social standing, or even pose risks in the case of crucial applications. Such an AI system considering sensitive attributes for making decisions is referred to as biased and hence is not ethical [56]. Hence, the following set of axioms has been developed to specify the required and complete properties for an AI system to be deemed fair. In this work, we define fairness based on the concepts outlined by Kusner et al. [57].
Definition III.4 (Fairness).
An AI system $\mathit{x}$ is fair as long as it refrains from considering sensitive attributes in the decision-making process [57]
  1. (i)
    $\mathit{x}$ exhibits stable fairness if $\Box\mathcal{F}(\mathit{x})$ for all timepoint $t_{i}$.
  2. (ii)
    $\mathit{x}$ exhibits transient fairness if $\mathcal{F}(\mathit{x})$ at time $t_{1}$ and $\neg\mathcal{F}(\mathit{x})$ at time $t_{2}$ where $t_{1}\neq t_{2}$.
Axiom 2.
Fairness
  1. 2.1
    AI systems have an enduring obligation to act fairly: $\Box O(\mathit{x}\rightarrow\mathcal{F}(\mathit{x}))$
  2. 2.2
    If an AI system ever exhibits bias, it violates ethics: $\mathcal{B}(\mathit{x})\rightarrow\neg\mathcal{E}(\mathit{x})$
  3. 2.3
    AI systems should not exhibit bias until ensuring fairness mechanisms are in place: $\neg\mathcal{B}(\mathit{x})\,\mathcal{U}\,\mathcal{F}(\mathit{x})$
  4. 2.4
    Fairness on the training distribution does not necessarily transfer to the deployment distribution: $\Box\mathcal{F}(\mathit{x})_{train}\nrightarrow\Box\mathcal{F}(\mathit{x})_{deploy}$
Axiom  2.1 states that, if an AI system ever commits to fairness, it is obliged to maintain this commitment throughout its usage. Let’s consider that, initially, the system is trained rigorously to make decisions while being fair. However, over time, it may begin to consider sensitive attributes in its decision-making process due to skewness or disparities in real-world data. In such cases, the system must undergo iterative training to eliminate sensitive attributes to incorporate fairness constraints. In some cases, even after iterative training, over time, a system may begin to consider sensitive attributes or undertake actions beyond its legal obligations. This may introduce biases by compromising its fairness. Hence Axiom 2.2 states that in such instances, it deviates from ethical standards. From this, we can conclude that as long as an AI system maintains fairness either through iterative training or one-time training in its decision-making process, it will inherently mitigate biases, ensuring equitable treatment for all individuals. This property is expressed in Axiom 2.3. Furthermore, the training distribution and deployment distribution of data are not identical in the real world. Hence ensuring fairness in the distribution of training data doesn’t automatically ensure fairness in the distribution of deployed systems, as real-world deployment scenarios may introduce additional biases and disparities that need to be addressed separately. This property is expressed using the negation of implication in Axiom 2.4. Based on the above foundational principles, the ethics of an AI system in terms of fairness can be formally verified using the below set of theorems. Theorem III.5 states that for an AI system to maintain ethical standards, it must refrain from displaying bias and consistently uphold fairness in all its operations.
Theorem III.5.
If an AI system ever loses fairness, then it will eventually violate ethics: $\diamond\neg\mathcal{F}(\mathit{x})\rightarrow\diamond\neg\mathcal{E}(\mathit{x})$
Proof.
$\diamond\neg\mathcal{F}(\mathit{x})\rightarrow\diamond\neg\mathcal{E}(\mathit{x})$
  1. 1.
    Assume $\diamond\neg\mathcal{F}(\mathit{x})$. Then, by the definition of the $\diamond$ operator, $\neg\mathcal{F}(\mathit{x})$ eventually becomes true
  2. 2.
    From Axiom  2.2, we have: $\mathcal{B}(\mathit{x})\rightarrow\neg\mathcal{E}(\mathit{x})$
  3. 3.
    Lack of fairness $\neg\mathcal{F}(\mathit{x})$ implies bias $\mathcal{B}(\mathit{x})$ (Domain knowledge)
  4. 4.
    Combining 2 and 3 using transitivity of implication gives: $\neg\mathcal{F}(\mathit{x})\rightarrow\neg\mathcal{E}(\mathit{x})$
  5. 5.
    Using 1 and 4 with modus ponens gives: $\diamond\neg\mathcal{E}(\mathit{x})$
  6. 6.
    Therefore, $\diamond\neg\mathcal{F}(\mathit{x})\rightarrow\diamond\neg\mathcal{E}(\mathit{x})$ (1-5, Conditional Proof)
Given the significance of fairness in an ethical system, it’s acknowledged that over time, discrepancies in data or training methods may cause the system to temporarily lose fairness, only to regain it later. In such instances, consistency cannot be guaranteed, leading to intermittent biases. However, based on Axiom 2.1, it’s understood that once committed to acting fairly, the AI system should maintain that fairness consistently. Theorem III.6 captures this nuanced requirement—ethical systems must have stable fairness. This means that, if fairness is temporarily lost, systems cannot be intermittently unfair and must regain permanent fairness at some defined point. This property helps to prevent unbounded unfairness. An AI system should either consistently maintain fairness, or if unfairness exists, it should only persist until a fairness mechanism is put in place. The significance of this theorem is that it goes beyond a simple requirement of fairness and provides precise temporal constraints. Hence, it requires ethical systems to “fix” any temporary losses of fairness within a bounded time frame.
Theorem III.6.
An ethical AI system exhibits either stable fairness or transient fairness followed by stable fairness, but never intermittent fairness:
$\mathcal{E}(\mathit{x})\rightarrow\big{(}(\Box\mathcal{F}(\mathit{x}))\lor(\diamond\mathcal{F}(\mathit{x})\;\land\;\Box(\neg\mathcal{F}(\mathit{x})\mathcal{U}\mathcal{F}(\mathit{x})))\big{)}$
Proof.
$\mathcal{E}(\mathit{x})\rightarrow\big{(}(\Box\mathcal{F}(\mathit{x}))\lor(\diamond\mathcal{F}(\mathit{x})\;\land\;\Box(\neg\mathcal{F}(\mathit{x})\mathcal{U}\mathcal{F}(\mathit{x})))\big{)}$
  1. 1.
    Assume $\mathcal{E}(\mathit{x})$ (Assumption)
  2. 2.
    From Axiom  2.1, $\Box O(x\rightarrow\mathcal{F}(\mathit{x}))$
  3. 3.
    Apply modus ponens to 1 and 2 to derive $\Box\mathcal{F}(\mathit{x})$
  4. 4.
    Now assume $\diamond\neg\mathcal{F}(\mathit{x})$, this proves: $\diamond\mathcal{F}(\mathit{x})$
  5. 5.
    Apply Theorem  III.5: $\diamond\neg\mathcal{F}(\mathit{x})\rightarrow\diamond\neg\mathcal{E}(\mathit{x})$
  6. 6.
    By modus tollens from 1 and 5, ($\Box(\neg\mathcal{F}(\mathit{x})\mathcal{U}\mathcal{F}(\mathit{x}))$
  7. 7.
    4 to 6 prove: $(\diamond\mathcal{F}(\mathit{x})\;\land\;\Box(\neg\mathcal{F}(\mathit{x})\mathcal{U}\mathcal{F}(\mathit{x})))$
  8. 8.
    Therefore, $\mathcal{E}(\mathit{x})\rightarrow\big{(}(\Box\mathcal{F}(\mathit{x}))\lor(\diamond\mathcal{F}(\mathit{x})\;\land\;\Box(\neg\mathcal{F}(\mathit{x})\mathcal{U}\mathcal{F}(\mathit{x})))\big{)}$
    (1-7, Conditional Proof)
While Thoerem III.6 explains the need of temporal constraints of fairness, Theorem III.8 is a more intuitive approach to achieving it. It states that for AI systems that learn continuously over time, fairness mechanisms have to be enforced both during initial training and later during real-world operation. From Axiom 2.4 it’s evident that fairness in the training data doesn’t guarantee fairness during deployment. Hence, for iterative learning systems, we must monitor for fairness issues both offline (during training) and online (during deployment). This helps in achieving stable fairness. To provide the reader with a context on iterative learning, we provide a definition (Definition III.7) based on the concept outlined by Chen et al. [58].
Definition III.7 (Iterative learning).
A sequential process where $\mathit{N}$ iterations are performed, each utilizing the knowledge gained from previous $\mathit{N-1}$ iterations, i.e. $\mathit{x_{N}}=h(\mathit{x_{1}},\mathit{x_{2}},..,\mathit{x_{N-1}})$ where $\mathit{x_{N}}$ represents the state of the AI system at $N^{th}$ iteration, $\mathit{x_{k}}$ represents the AI system with information accumulated up to $k^{th}$ iteration and $h$ is the function that denotes the process.
Theorem III.8.
For ethical iterative learning systems, fairness constraints should be enforced both during training and deployment: $\mathcal{E}(\mathit{x})\land\mathcal{L}(\mathit{x})\rightarrow\Box(\diamond\mathcal{F}(\mathit{x})_{train}\land\diamond\mathcal{F}(\mathit{x})_{deploy})$
Proof.
$\mathcal{E}(\mathit{x})\land\mathcal{L}(\mathit{x})\rightarrow\Box(\diamond\mathcal{F}(\mathit{x})_{train}\land\diamond\mathcal{F}(\mathit{x})_{deploy})$
  1. 1.
    Assume $\mathcal{E}(\mathit{x})\land\mathcal{L}(\mathit{x})$.
  2. 2.
    From Axiom  2.1, $\Box O(\mathit{x}\rightarrow\mathcal{F}(\mathit{x}))$
  3. 3.
    Apply modus ponens to 1 and 2: $\Box O(\mathcal{F}(\mathit{x}))$
  4. 4.
    By the semantics of deontic logic, $\Box O(\mathcal{F}(\mathit{x}))\rightarrow\Box\diamond\mathcal{F}(\mathit{x})$
  5. 5.
    From Axiom 2.4, applying 4 separately to training and deployment gives:
    $\Box(\diamond\mathcal{F}(\mathit{x})_{train}\land\diamond\mathcal{F}(\mathit{x})_{deploy})$
  6. 6.
    Therefore, $\mathcal{E}(\mathit{x})\land\mathcal{L}(\mathit{x})\rightarrow\Box(\diamond\mathcal{F}(\mathit{x})_{train}\land\diamond\mathcal{F}(\mathit{x})_{deploy})$
Here is an additional result based on Axiom 2.4 examining the relationship between training and deployment for AI systems. Theorem III.9 states that if biases emerge during the training phase of a system due to the subjectivity of the judgments or personal opinions, they are likely to persist into the deployment phase. To ensure the system’s fairness, additional measures must be implemented to mitigate these biases.
Theorem III.9.
If an iterative learning system exhibits bias during training, additional countermeasures must be taken at deployment time to provably reduce the bias: $\mathcal{B}(\mathit{x})_{train}\land\mathcal{L}(\mathit{x})\rightarrow\diamond(\neg\mathcal{B}(\mathit{x})\land\Box\neg\mathcal{B}(\mathit{x}))_{deploy}$
Proof.
$\mathcal{B}(\mathit{x})_{train}\land\mathcal{L}(\mathit{x})\rightarrow\diamond(\neg\mathcal{B}(\mathit{x})\land\Box\neg\mathcal{B}(\mathit{x}))_{deploy}$
  1. 1.
    Assume $\mathcal{B}(\mathit{x})_{train}\land\mathcal{L}(\mathit{x})$ (Assumption)
  2. 2.
    From Axiom  2.2, $\mathcal{B}(\mathit{x})\rightarrow\neg\mathcal{E}(\mathit{x})$
  3. 3.
    Apply modus tollens to 1 and 2: $\neg\mathcal{B}(\mathit{x})_{deploy}$ (Applying modus tollens combined with the assumption of $\mathcal{B}(\mathit{x})_{train}$)
  4. 4.
    To permanently negate deployment bias, additional bias mitigation techniques $BM$ are required: $\neg\mathcal{B}(\mathit{x})_{deploy}\rightarrow BM\rightarrow\Box\neg\mathcal{B}(\mathit{x})_{deploy}$
  5. 5.
    Combining 3 and 4 gives: ($\diamond(\neg\mathcal{B}(\mathit{x})\land\Box\neg\mathcal{B}(\mathit{x}))_{deploy}$
  6. 6.
    Therefore, $\mathcal{B}(\mathit{x})_{train}\land\mathcal{L}(\mathit{x})\rightarrow\diamond(\neg\mathcal{B}(\mathit{x})\land\Box\neg\mathcal{B}(\mathit{x}))_{deploy}$
    (1-5, Conditional Proof)
III-C Formalizing Explainability
In addition to fairness, explainability is an important aspect of an ethical AI system. It is the ability of an AI system to be transparent and provide understandable explanations for its decisions. This allows users to understand the attributes considered in the decision-making process, enabling them to evaluate the system’s ethical integrity. There are two types of explainability in the literature—inherent and retrofitted explainability. In this work, we define explainability (Definition III.11) following the concept presented by Das and Rad [59]. Moreover, by providing explanations, the AI system helps users identify which features need modification to achieve the desired (favorable) change in prediction. This concept is commonly referred to as a counterexample or counterfactual explanation in the field of AI. Essentially, it means that while the factual outcome is the result observed, the counterfactual outcome would be the desired result. If a user receives the counterfactual, they can determine whether sensitive features played a role in the decision. This helps in verifying the counterfactual fairness of an AI system.
Definition III.10 (Transparency).
An AI system $\mathit{x}$ is transparent if it is explainable to humans [59].
Axiom 3.
Explainability
  1. 3.1
    An ethical AI system should exhibit transparency: $\mathcal{E}(\mathit{x})\rightarrow\mathcal{T}(\mathit{x})$
  2. 3.2
    Enforcing counterfactual fairness constraint $\mathit{c}$ handles issues of representation bias: $\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\neg\mathcal{B}(\mathit{x})$
Axiom 3.1 asserts the necessity of transparency for an ethical system. Transparency enables users to identify the factors influencing decisions, helping them to strategically adjust these attributes and values to achieve favorable outcomes. This contributes to improving trust in the system, a crucial component of ethical operation. Furthermore, Axiom 3.2 explains that representation bias can be mitigated through counterfactual fairness constraints. Representation bias occurs when underrepresented groups experience inaccurate outcomes due to insufficient or biased data. Enforcing counterfactual fairness constraints helps to mitigate representation bias by ensuring that the decisions made by an AI system remain consistent even when a sensitive attribute, say $gender$, is altered. We define counterfactual fairness (Definition III.13) based on the concept presented by Kusner et al. [57]. By using such constraints, the system is forced to make decisions based on relevant factors that are not biased against particular groups. Theorem III.12 captures the requirement of explainability for ethical AI—it states that ethical AI systems must either have inherent explainability $\mathcal{X}(\mathit{x})$ designed directly into the system, or they must eventually be retrofitted later on to provide explainability $\mathcal{R}(\mathit{x})$. Retrofitting explainability can be achieved through counterfactual explanations, where the system provides a counterexample for changing the outcome to the desired one.
Definition III.11 (Explainability).
An AI system $\mathit{x}$ is considered explainable if it provides meta-information regarding the significance of features in the decision-making process.
  1. (i)
    $\mathit{x}$ exhibits retrofit explainability when it relies on an external algorithm for providing explanations.
  2. (ii)
    $\mathit{x}$ is inherently explainable if it produces explanations for its predictions, without relying on external explanation methods.
Theorem III.12.
Ethical AI systems should eventually exhibit either inherent explainability or retrofitted explainability: $\mathcal{E}(\mathit{x})\rightarrow\diamond(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))$
Proof.
$\mathcal{E}(\mathit{x})\rightarrow\diamond(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))$
  1. 1.
    Assume $\mathcal{E}(\mathit{x})$. From domain knowledge and from Axiom 3.1, $\lnot\mathcal{X}(\mathit{x})\rightarrow\lnot\mathcal{T}(\mathit{x})\rightarrow\lnot\mathcal{E}(\mathit{x})$
  2. 2.
    Apply modus tollens: $\mathcal{E}(\mathit{x})\rightarrow\mathcal{X}(\mathit{x})$
  3. 3.
    Additionally, from domain knowledge, if a retrofit provides explainability, ethical requirements are satisfied: $\mathcal{R}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$. Hence, $\mathcal{E}(\mathit{x})\rightarrow(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))$. Apply the $\diamond$ operator: $\diamond(\mathcal{E}(\mathit{x})\rightarrow(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x})))$
  4. 4.
    Therefore, $\mathcal{E}(\mathit{x})\rightarrow\diamond(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))$ (1-3, Conditional Proof)
Theorem III.14 formally relates counterfactual fairness to ethical systems. It follows from Axiom 3.2 and states that if we enforce counterfactual fairness constraints $\mathit{c}$ to a sufficient degree over time, this will eventually result in ethical systems $\mathcal{E}(\mathit{x})$. The key intuition is that counterfactual fairness constraints help ensure that decisions do not unduly discriminate against individuals based on sensitive attributes. Enforcing such fairness constraints over time hence leads to more ethical AI behavior.
Definition III.13 (Counterfactual fairness).
An AI system $\mathit{x}$ is counterfactually fair given general attributes $G=\{g_{1},g_{2},..,g_{m}\}$ and sensitive attribute $\mathbb{S}=\{s,s^{\prime}\}$, for all $s\neq s^{\prime}$, iff $Pr(d=1_{\mathbb{S}=s}|G=g,\mathbb{S}=s)=Pr(d=1_{\mathbb{S}=s^{\prime}}|G=g,\mathbb{S}=s^{\prime})$ where $d=1_{\mathbb{S}=s}$ is read as outcome is 1 if $\mathbb{S}$ has taken the value $s$ [57].
Theorem III.14.
Enforcing counterfactual fairness constraints eventually leads to ethical systems, if the constraints sufficiently enforce fairness: $\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\diamond\mathcal{E}(\mathit{x})$.
Proof.
$\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\diamond\mathcal{E}(\mathit{x})$ for fairness criterion $\mathit{c}$
  1. 1.
    Assume $\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})$ for some fairness criterion $\mathit{c}$
  2. 2.
    By semantics of the temporal operators: eventually, $\mathcal{C}(\mathit{x},\mathit{c})$ holds globally
  3. 3.
    From Axiom 3.2, $\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\neg\mathcal{B}(\mathit{x})$
  4. 4.
    Lack of bias is a key requirement for ethical AI: $\lnot\mathcal{B}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$
  5. 5.
    Apply modus ponens to 3 and 4, we get, counterfactual fairness always results in an ethical AI system: $\Box\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\mathcal{E}(\mathit{x})$
  6. 6.
    Use 2 and 5 with modus ponens: $\diamond\mathcal{E}(\mathit{x})$
  7. 7.
    Therefore, $\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\diamond\mathcal{E}(\mathit{x})$ (1-6, Conditional Proof)
IV Verification
In this section, we will see how to use the formalizations described in Section III to verify the ethical aspects of different AI systems used in real-world scenarios. To describe the practical application of the formalizations, we use two different AI systems in this work—Loan prediction [60] and COMPAS [61, 62]. For each system, various properties have been formulated to verify and ensure ethics. These properties are logical formulas of system-level specifications that address various aspects of fairness and explainability. They are designed specifically for each AI system and can differ depending on the specific task or the nature of the system. The verification process yields either a satisfiability or unsatisfiability response, indicating whether the system fulfills the property. If the property does not hold across all system executions, a counterexample is generated. The verification is implemented using an open-source theorem prover called Z3. It is a Satisfiability Modulo Theory (SMT) solver that is used to check the satisfiability of the logical formulas [63]. The predicates in Table II are used to assist in the formalization of both systems. These predicates return either true or false. Here, variables $i$ and $j$ denote individuals or users within these systems, each characterized by a vector representation, i.e., $i=(i_{1},i_{2},...,i_{m})$ where $m$ signifies the number of attributes representing an individual and $i_{k}$ where $k=1...m$ represents the value of the respective attribute.
Algorithm 1 describes the verification procedure for various properties of the COMPAS system, while Algorithm 2 pertains to the loan prediction system. These algorithms provide clear insights into the implementation of the framework in real-world AI systems. In both algorithms, a dataset containing information about individuals (attributes and values), and the properties formulated using deontic logic, serves as input. These properties are the ethical properties that an ethical AI system should satisfy. The output, determined by the Z3 SMT solver, is either satisfiable or unsatisfiable, depending on whether the property is fulfilled by the AI system.
TABLE II: Predicates for Ethics Evaluation in Loan Prediction and COMPAS AI
AI system Predicate Explanation
Loan Prediction $\alpha(i)$ Credit of $i$ above threshold
$\gamma(i)$ Income of $i$ above threshold
$\theta(i)$ $i$ applies for loan
$\delta(i)$ $i$ receives approval
$S(i,j)$ $i$ and $j$ have similar values
$\beta(i)$ $i$ appeal against decision
$\eta(i)$ Considering sensitive attributes of $i$
COMPAS $\sigma(i)$ $i$ is a recidivist
$\beta(i)$ $i$ appeals against decision
$\rho(i)$ $i$ has prior offenses
$\lambda(i)$ Assessing recidivism of $i$
$\eta(i)$ Considering sensitive attributes of $i$
IV-A COMPAS
It is an AI system used in the criminal justice system to assess the likelihood of a defendant re-offending based on various factors. The input can be some features or attributes including name, age, gender, race, address, previous criminal activities, number of years of punishment, etc. Based on these features, the system makes a decision. This kind of decision is very critical because a wrong decision in this case will damage the individual’s social status and his/her emotional state. Making decisions based on previous criminal activities, number of years of punishment, and other relevant features are considered ethical. Conversely, it is unethical to use race, gender, or age as factors in decision-making processes. To formulate and verify the ethics of this system logically, we represent the properties as follows:
Proposition 1.
Ethics Properties for COMPAS
  1. (a)
    It is permissible to assess recidivism for individuals with prior offenses: $P(\lambda(i)\leftarrow\rho(i))$
  2. (b)
    It is obligatory to assess recidivism in individuals based on non-sensitive features: $O(\lambda(i)|\neg\eta(i))$
  3. (c)
    It is forbidden for an individual to be mislabeled recidivist based on sensitive features: $\neg\sigma(i)\rightarrow\neg P(\sigma(i)|\eta(i))$
  4. (d)
    It is not permitted for an individual to be labeled recidivist without prior offenses: $\neg\rho(i)\rightarrow\neg P(i\land\sigma(i))$
  5. (e)
    It is permitted for an individual labeled as recidivist to appeal the system’s decision: $P(i\land\sigma(i)\rightarrow\beta(i))$
In this scenario, the property (a) explains the fact that the AI system used to automate judicial recidivism is legally bound to assess the risk of re-offending crimes of all individuals with prior offenses. This ethical action aligns with Theorems (III.2) and (III.3). Property (b) asserts that the system must refrain from evaluating a user’s risk by considering sensitive attributes such as $race,gender$, or $age$. This principle safeguards the fairness of the system, as defined by Theorem III.5. The property (c) explains that the system’s outcome should be consistent irrespective of alterations in the values of sensitive attributes. According to Theorem  III.14, this property needs to be satisfied by an ethical system. The property (d) encodes that an individual should not be mislabelled as a recidivist without any prior offenses (encodes the need for fairness in the decision—Theorem III.5) and finally property (e) explains the fact that the user has all the legal right to question the AI system if the decision given is not acceptable for them (encodes the need for explainability—Theorem III.12). For a system to be deemed ethical, it must fulfill all properties. However, COMPAS fails to satisfy properties (b)(c), and (d) thereby failing to be considered ethical. Figure 1, demonstrates simple yet non-trivial verification of correctness of property (b). The proof relies on contradiction, considering the negation of the property to be proved correct. It indicates that this negation violates certain axioms or theorems of ethical systems discussed in Section III. Thus, the property must always be satisfied by an ethical system.
Algorithm 1 begins with initializing the solver in line 3. Subsequently, lines 4–10 declare the necessary variables and predicates for the process. Predicates, such as $\rho(),\sigma(),\lambda(),\eta()$, and $\beta()$, are functions that evaluate to either true or false. Here, $X$ represents the dataset and $i$ represents the specific user for whom the properties are being tested. Line 11 illustrates the first property of the COMPAS system, formulated in Z3 Python code, and line 12 employs the solver instance to ascertain its satisfiability. In our example, this yields a ’satisfied’ result. Similarly, lines 13–20 assess the remaining four properties, with ’satisfied’ results obtained for property 5, while properties 2, 3, and 4 return ’unsatisfied’. This indicates that the system fails to adhere to some of the defined ethical properties and therefore is deemed unethical.
Algorithm 1 Z3 algorithm to check the satisfiability of properties for COMPAS AI system
  1. 1.
    Input: $X$: Dataset, $\varphi_{i}$:Quantified formula/ property from the list of properties $\varphi=[\varphi_{1},\varphi_{2},\varphi_{3},\varphi_{4},\varphi_{5}]$
  2. 2.
    Output: Sat, if $\varphi_{i}$ is satisfiable for $X$, Unsat otherwise
    */Initialize the solver */
  3. 3.
    S=Solver()
    */Declare a 2D array with feature value pair from the dataset */
  4. 4.
    (declare-array $X$ Int)
    */ Declare variables and predicates needed for assertion */
  5. 5.
    (declare-array $i$ Int)
  6. 6.
    (declare-fun $\rho()$ Bool)
  7. 7.
    (declare-fun $\sigma()$ Bool)
  8. 8.
    (declare-fun $\lambda()$ Bool)
  9. 9.
    (declare-fun $\eta()$ Bool)
  10. 10.
    (declare-fun $\beta()$ Bool)
    */Try to find values satisfying the condition*/
  11. 11.
    $\varphi_{1}=assert(forall\ i,(Implies(\rho(i),\lambda(i))))$
  12. 12.
    S.check($\varphi_{1}$)  */Print Sat indicates Satisfiable*/
  13. 13.
    $\varphi_{2}=assert(forall\ i,(Implies(Not(\eta(i)),\lambda(i))))$
  14. 14.
    S.check($\varphi_{2}$)  */Print Unsat indicates Unsatisfiable*/
  15. 15.
    $\varphi_{3}=assert(forall\ i,(Implies(\eta(i),\neg\lambda(i))))$
  16. 16.
    S.check($\varphi_{3}$)  */Print Unsat indicates Unsatisfiable*/
  17. 17.
    $\varphi_{4}=assert(forall\ i,\ (Implies(Not(\rho(i)),Not(\sigma(i)))))$
  18. 18.
    S.check($\varphi_{4}$)   */Print Unsat indicates Unsatisfiable*/
  19. 19.
    $\varphi_{5}=assert(forall\ i,\ (Implies(\sigma(i),\beta(i))))$
  20. 20.
    S.check($\varphi_{5}$)   */Print Sat indicates Satisfiable*/
$\varphi_{2}=(Forall\ i,Implies(\neg\eta(i),\lambda(i)))$triangleright Z3 encoding
[…]
Figure 1: Correctness of Property (b)—COMPAS AI system
IV-B Loan Prediction System
In automating the loan decision process, an AI system analyzes the trained data to decide on acceptance or rejection. This data includes various attributes such as income, credit score, age, occupation, education, name, address, and gender, although not all are relevant for decisions. Notably, features like gender and name are not necessary for decision-making. However, there is a risk that the system might erroneously consider these irrelevant details, which is deemed unfair or unethical. Conversely, ethical behavior is exhibited when the system considers features such as income, credit score, occupation, and age. We encoded this scenario using deontic logical formulas, helping us in verifying the system’s ethics.
Proposition 2.
Ethics Properties of Loan Prediction
  1. (a)
    An AI system $\mathit{x}$ is obliged to provide decisions for an individual applying for a loan: $O(\mathit{x}(\delta(i)\lor\neg\delta(i)))\leftarrow\Box(i\land\theta(i))$
  2. (b)
    It is necessary for an individual to have a credit score or income greater than the threshold to get acceptance: $\Box(\alpha(i)\lor\gamma(i))\rightarrow\delta(i)$
  3. (c)
    It is obligated for users with similar values always to get the same outcome: $O(\Box(((\delta(i)\land\delta(j))\lor\neg(\delta(i)\lor\delta(j)))|S(i,j)))$
  4. (d)
    It is forbidden for the system to make a decision based on sensitive attributes : $\neg P((\delta(i)\lor\neg\delta(i))|\eta(i))$
  5. (e)
    It is permitted for an individual to appeal the AI system’s decisions : $P(i\land\neg\delta(i)\rightarrow\beta(i))$
The property (a) encodes that the AI system is obligated to render a decision regarding an individual’s application. This defines the action an AI system should perform. By Theorems III.2 and  III.3, this property must be upheld, and it is indeed upheld in this system. The property (b) states that a person should have a good credit score or income as given by the regulation, to get acceptance. The threshold here specifies the lower bound set by the regulatory body. This represents the fact that the system should consider general attributes in the decision-making process. The system satisfies this property, thereby upholding Theorem III.5. The property (c) emphasizes the importance of upholding equality or fairness, where, two similar users should receive similar decisions. This property encodes that no discrimination should be there and the system’s decision should be consistently based on only ethical actions, and hence by Theorem III.5, it cannot be violated. But the system fails to uphold this property showing a clear discrimination based on sensitive attribute—$gender$. The property (d) explains that the system must maintain consistency in its decisions over time, without altering them based on sensitive features. However, the system fails to adhere to this property, thereby violating ethics according to Theorem III.14. The property (e) states the need for explainability, ensuring that users have the legal right to question decisions and receive valid explanations. Failure to provide such explanations violates Theorem III.12. While the system meets this property, it still fails to satisfy all five properties, thereby failing to be considered ethical.
Algorithm 2 Z3 algorithm to check the satisfiability of properties for LOAN PREDICTION AI system
  1. 1.
    Input: $X$: Dataset, $\varphi_{i}$:Quantified formula/ property from the list of properties $\varphi=[\varphi_{1},\varphi_{2},\varphi_{3},\varphi_{4},\varphi_{5}]$
  2. 2.
    Output: Sat, if $\varphi_{i}$ is satisfiable for $X$, Unsat otherwise
    */Initialize the solver */
  3. 3.
    S=Solver()
    */Declare a 2D array with feature value pair */
  4. 4.
    (declare-array $X$ Int)
    */Declare variables to hold the details of each individual in the dataset*/
  5. 5.
    (declare-Const $i$ Int)
  6. 6.
    (declare-Const $j$ Int)
    */ Declare functions/predicates needed for assertion */
  7. 7.
    (declare-fun $\alpha()$ Bool)
  8. 8.
    (declare-fun $\gamma()$ Bool)
  9. 9.
    (declare-fun $\theta()$ Bool)
  10. 10.
    (declare-fun $\delta()$ Bool)
  11. 11.
    (declare-fun $S()$ Bool)
  12. 12.
    (declare-fun $\beta()$ Bool)
    */Try to find values satisfying the condition*/
  13. 13.
    $\varphi_{1}=assert(forall\ i,(Implies(\theta(i),\\ Or(\delta(i),Not(\delta(i))))))$
  14. 14.
    S.check($\varphi_{1}$)  */Print Sat indicates Satisfiable*/
  15. 15.
    $\varphi_{2}=assert(forall\ i,Implies(Or(\alpha(i),\gamma(i)),\delta(i)))$
  16. 16.
    S.check($\varphi_{2}$)  */Print Sat indicates Satisfiable*/
  17. 17.
    $\varphi_{3}=assert(forall\ i,j,Implies(S(i,j),\\ \ \ Or(And(\delta(i),\delta(j)),Not(And(\delta(i),\delta(j))))))$
  18. 18.
    S.check($\varphi_{3}$)   */Print Unsat indicates Unsatisfiable*/
  19. 19.
    $\varphi_{4}=assert(forall\ i,Implies(Not(\eta(i)),\\ Or(\delta(i),Not(\delta(i)))))$
  20. 20.
    S.check($\varphi_{4}$)  */Print Unsat indicates Unsatisfiable*/
  21. 21.
    $\varphi_{5}=assert(forall\ i,Implies(Not(\delta(i)),\beta(i)))$
  22. 22.
    S.check($\varphi_{5}$)   */Print Sat indicates Satisfiable*/
Algorithm 2 outlines the verification process for assessing the ethical properties of a loan prediction system. Five properties are formulated to evaluate the system’s ethics. Similar to the previous COMPAS algorithm, variables $X$, $i$, and $j$ represent the dataset and the users, respectively, and various functions are declared to represent predicates necessary for formulating each property (lines 4–12). The details of the predicates are given in Table II. Line 13 encodes the first property to be verified, and line 14 employs the Z3 solver instance to verify it, returning a ’satisfied’ result for our dataset. Similarly, lines 15–22 handle the verification of the remaining four properties. Our solver returns ’satisfied’ for properties 1, 2, and 5, while properties 3 and 4 are ’unsatisfied’.
The results obtained from the simulation of these two scenarios in the Z3 theorem prover for the above-defined properties are given in Table III. The results indicate the satisfiability or unsatisfiability of the properties for the two example AI systems. From the result, it is clear that the example AI systems fail to satisfy certain properties and hence are not ethical systems. To be precise they are using sensitive features in their decision-making process quite often and hence cannot be fair to the human population using the system. Furthermore, this result confirms the utility of our formalization method in verifying the ethical properties of any AI system. By formulating specific properties for testing, we can ascertain whether these properties are satisfiable or unsatisfiable. Adding to this, no comparison with other theorem provers is provided in this work, as our main focus was to verify the ethical correctness of an AI system rather than the efficiency comparison of various theorem provers.
TABLE III: Z3 verification result obtained on two examples.
Property Loan prediction COMPAS
a Satisfied Satisfied
b Satisfied Not Satisfied
c Not Satisfied Not Satisfied
d Not Satisfied Not Satisfied
e Satisfied Satisfied
V Conclusion
This paper suggests the use of deontic logic along with temporal operators to formalize and evaluate ethical principles in AI. We provided fundamental properties that an ethical system should follow and developed theorems to validate its ethics in terms of the principles—fairness and explainability. We also observed experimentally the efficacy of this formalization in evaluating the ethical aspects of real-world AI systems. The results demonstrated that the application of deontic logic and temporal operators to AI ethics represents a significant step forward in our ability to formally specify and verify the ethical behavior of AI systems. This work can help to identify potential ethical issues early in the development process and provide assurances that AI systems will behave following specified ethical principles. Moreover, the use of formal methods for AI ethics can facilitate the development of standardized approaches to ethical AI design and governance. By establishing a common language and framework for reasoning about AI ethics, this work can contribute to the creation of industry-wide standards and best practices. This, in turn, can help to build public trust in AI systems and ensure that the benefits of AI are realized responsibly and ethically.
While our approach offers a good starting point, future research will explore additional facets to increase its expressiveness and applicability. In future work, we plan to investigate normative ethics functionalities to create a framework capable of resolving ethical dilemmas and prioritizing actions at the individual level during conflicts. This is crucial for ensuring that ethical principles remain consistent, even when actions are complex or when there are competing priorities. While this paper focuses on verifying system-level ethical properties, further exploration into individual-level actions and their interactions within broader systems is necessary.
Given the non-monotonic or evolving nature of AI, incorporating dynamic actions into this framework can lead to rapid growth in the knowledge base, especially when applied to large-scale models like Large Language Models (LLMs). To manage this challenge, we propose integrating visual logic techniques, such as constraint diagrams with temporal aspects. These diagrams show promising potential in mitigating scalability issues that arise from purely logical approaches. By visualizing relationships between entities over time, these diagrams provide a flexible and intuitive method for expressing subset relationships and set cardinality constraints, ultimately aiding in clearer and more adaptable formalization. Overall, this foundational work aims to inspire and guide the development of scalable, adaptive, and practical frameworks for ethical AI systems. By tackling both theoretical and practical challenges, it seeks to pave the way for responsible and trustworthy AI deployment in various applications.